OpenStudy (anonymous):

The halflife of a radioactive radium is 1620 years. What percent of a present amount of radioactive radium will remain after 100 years?

OpenStudy (anonymous):

every 1620 years, the amount decreases by 1/2 so 2^-(t/1620)

OpenStudy (anonymous):

t would be the amount of years, so when t = 0, then 2^-(t/1620)= 1 when t=1620, then 2^-(t/1620)= 1/2, and so on and so forth so now solve for when t=100